Joe has a collection of nickels and dimes that is worth $5.65. If the number of dimes were doubled and the number of nickels were increased by 8, the value of the coins would be $10.45. How many dimes does he have?
Answer by mangopeeler07(462)

d=number of dimes
n=number of nickels
coefficient= value of the coin

value of dime(number of dimes)+value of nickel(number of nickels)=total value
10(d)+5(n)=565

10(2d)+5(n+8)=1045------------If the number of dimes were doubled and the number of nickels were increased by 8, the value of the coins would be $10.45.

Distribute
20d+5n+40=1045

Subtract 40 from both sides
20d+5n=1005
and
10d+5n=565 both have 5n, so subtract them from each other

10d=440

Divide both sides by 10
d=44

Number of dimes=44